OpenStudy (anonymous):

The radius of a sphere is increasing at a rate of 4 mm/s. How fast is the volume increasing when the diameter is 40 mm?

OpenStudy (perl):

V = 4/3*Pi*R^3 dV/dt = 4/3 * Pi * 3*R^2*dR/dt dV/dt = 4*Pi*R^2*dR/dt when diameter is 40 , Radius = 20. and dR/dt = 4. (make sure units are consistent) dV/dt = 4*Pi*(20)^2 * 4

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

hmm..